OpenStudy (anonymous):

What is the exact value of the expression 6the square root of y to the eighth power. − the square root of 9 times y to the fourth power. + 3the square root of y to the eighth power. − the square root of y to the fourth power.? Simplify if possible. y4 9y4 − 4y2 3y4 − 10y2 y2

OpenStudy (anonymous):

simplify it? i can barely read it

OpenStudy (anonymous):

6the square root of y to the eighth power. − the square root of 9 times y to the fourth power. + 3the square root of y to the eighth power. − the square root of y to the fourth power.

OpenStudy (anonymous):

\[\large 6\sqrt{y^8}-9y^4+3\sqrt{y^8}\]

OpenStudy (anonymous):

\[\large 6\sqrt{y^8}-9y^4+3\sqrt{y^8}-\sqrt{y^4}\]

OpenStudy (anonymous):

second one close?

OpenStudy (anonymous):

\[\large 6\sqrt{y^8}-9\sqrt{y^4}+3\sqrt{y^8}-\sqrt{y^4}\]

OpenStudy (anonymous):

how about that one?

OpenStudy (anonymous):

seccood one 9 has a square root sign

OpenStudy (anonymous):

\[\large 6\sqrt{y^8}-\sqrt{9y^4}+3\sqrt{y^8}-\sqrt{y^4}\]

OpenStudy (anonymous):

got it on the third try how many times does 2 go in to 8 ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

really it is that simple half of 8 is 4 so you get \[\large 6\sqrt{y^8}-\sqrt{9y^4}+3\sqrt{y^8}-\sqrt{y^4}\] \[=\large 6y^4-\sqrt{9y^4}+3y^4-\sqrt{y^4}\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

then half of 4 is 2 giving \[=\large 6y^4-\sqrt{9y^4}+3y^4-\sqrt{y^4}\] \[==\large 6y^4-3y^2+3y^4-y^2\]

OpenStudy (anonymous):

then combine like terms you good from there or no?

OpenStudy (anonymous):

b right

OpenStudy (anonymous):

right